DISSOCIATION ENERGIES OF SIX NO_{2} ISOTOPOLOGUES

\underline{REMY JOST}, M. JOYEUX and D. SUGNY, Laboratoire de Spectrométrie Physique,

Université J. Fourier–Grenoble I, B.P. 87 – 38402 Saint Martin d’Hères Cedex, FRANCE; G. MICHALSKI and M. THIEMENS, University of California, San Diego La Jolla Ca USA.

We have measured the dissociation threshold energy, D_{0}, (NO_{2}+hν → NO(^{2}Π_{1/2})+O(^{3}P_{2})) of the six NO_{2} isotopologues made with ^{14}N or ^{15}N and ^{16}O or ^{18}O isotopes. These NO_{2} isotopologues are cooled in a Helium supersonic jet at T_{rot} ≈ 2K. For each isotopologue, the very dense set of bound N = 1 K = 0 rovibronic eigenstates is readily observed by LIF up to D_{0}. Above D_{0}, the LIF signal disappear abruptly, within ±0.03cm^{−1} which is the average spacing between observed R_{0} lines just below D_{0}. Note that resonances (lifetime ≈ 10^{−10} sec.) located above D_{0} can be observed in absorption (by CRDS) but no fluorescence can be detected from these. The six measured D_{0} range from 25128.56cm^{−1} for ^{16}O^{14}N^{16}O, noted (646), to 25171.80cm^{−1} for (858). At the B.O. approximation, these six D_{0} should have a common D_{e}. The shifts between these six D_{0} are due to the ZPE shifts of NO_{2} and NO. We have used and check the following relation: D_{0}(^{x}O^{y}N^{z}O) = D_{e}(NO_{2})+ZPE (^{y}N^{z}O) − ZP E(^{x}O^{y}N^{z}O) The ZPEs of the various NO and NO_{2} isotopologues have been determined from Dunham parameters and, for NO_{2}, also by Canonical Perturbation Theory (CPT) using two PESs of NO_{2}. The NO_{2} ZPE isotopologue shifts are estimated to be within 0.5cm^{−1}. The uncertainties on ZPE of NO are significantly smaller. The six values of D_{e} are located within 0.5cm^{−1} around 26051.17cm^{−1}, in agreement with the ZPE uncertainties.